# Hypercube Policy Regularization Framework for Offline Reinforcement Learning

### v1

@@ 4.3 Algorithms
Input: Initial policy parameter ${\it\theta}$; Target policy parameter $\theta^{-}$; min-batch size $k$; Q-function parameters $\phi_{1}$,$\phi_{2}$; Target Q-function parameters $\phi_{1}^{-}$, $\phi_{2}^{-}$; offline dataset ${\it D}$; actions obtained by hypercube policy regularization ${a^{\prime}}$
@@
[…]
3:        Sample a mini-batch of $k$ state-action pairs $\left\{(s_{t},a_{t}),t=1,..,k\right\}$ from ${D}$
4:        Sample $a_{t+1}\sim\pi_{\theta}(a_{t+1}|s_{t+1})$
5:        Update $Q_{{\phi_{1}}}$ and $Q_{{\phi_{2}}}$ by $\begin{array}[]{l}E_{({s_{t}},a_{t},s_{t+1})\sim D,a_{t+1}\sim{\pi_{\theta}}}[(r({s_{t}},{a_{t}})\\ +\gamma\min Q_{{\phi_{i}}}({s_{t+1}},{a_{t+1}})-Q_{\phi_{i}}({s_{t}},{a_{t}}))^{2}]\end{array}$
6:        ##\## Policy learning
7:        Compared $\mathop{\min}\limits_{i=1,2}{Q_{{\phi_{i}}}}(s,a)$ and $\mathop{\min}\limits_{i=1,2}{Q_{{\phi_{i}}}}(s,a^{\prime})$ to Update ${a^{\prime}}$
8:        Update policy by Equation(15)
9:        # Update target network
10:        Update policy target by ${\theta^{-}}=\rho{\theta^{-}}+(1-\rho)\theta$
11:        Update Q target by: $\phi_{i}^{-}=\rho\phi_{i}^{-}+(1-\rho)\phi_{i}$
12:     end while
13:  end while
This section outlines how the hypercube policy regularization framework can be integrated with offline RL. The hypercube constraint framework can be directly applied to policy regularization algorithms such as TD3-BC and Diffusion-QL.
Taking TD3-BC as an example, our algorithm made only two straightforward modifications to TD3-BC. Firstly, a hypercube was constructed using Equation(3) and employed to store the optimal action within the hypercube. Each time the state $s$ corresponding to the hypercube is selected, a comparison is made between $Q(s,a)$ and $Q(s,a^{\prime})$ to determine whether action $a$ should be replaced within the hypercube. Secondly, in policy updates, the action $a_{h}$ from the hypercube is employed in place of $a$ from the static dataset for policy regularization. The policy update equation is given by:
$\arg\min L(\theta)={L_{c}}(\theta)+{L_{q}}(\theta)$ (15)
Where, $\theta$ is the policy parameter, and $L_{c}$ is the policy regularization loss function.
@@
$E_{(s,a)\sim D}[{(a-\pi_{\theta}(s))^{2}}]$ (16)
Unlike the $L_{c}$ above, the $L_{c}$ is:
$E_{(s,a)\sim D}[{(a_{h}-\pi_{\theta}(s))^{2}}]$ (17)
Where $a_{h}$ represents the action within the static dataset that has the highest Q within the hypercube. Given that $Q(s,a_{h})\geq Q(s,a)$ and Theorem 1, it follows that the actions $a_{h}$ within the hypercube will produce results that are either equal to or better than those produced by $a$ within the static dataset. Therefore, the utilization of Equation(3) for behavior cloning can lead to a more effective policy.
@@
$L_{q}=\phi E_{(s_{i},a_{i})\sim D}[\sum_{i=0}^{T-1}Q(s_{i},a_{i})]/E_{(s_{i},a_{i})\sim D}[\sum_{i=0}^{T-1}\left|Q(s_{i},a_{i})\right|]$ (18)
@@
Here $\gamma=\eta/{E_{(s_{i},a_{i})\sim D}}[\sum_{i=0}^{T-1}|{Q_{\phi}}(s,a)|]$, $\eta$ is a hyperparameter.
Env IQL CQL TD3-BC Diffusion-QL TD3-BC-C Diffusion-C
half-r $11.0{\pm}2.5$ $\textbf{28.3}{\pm}\textbf{0.5}$ $11.7{\pm}0.7$ $21.0{\pm}0.6$ $26.5{\pm}1.6$ $23.8{\pm}0.6$
hp-r $7.8{\pm}0.4$ $16.4{\pm}14.5$ $8.5{\pm}0.2$ $8.1{\pm}0.1$ $\textbf{31.3}{\pm}\textbf{0.0}$ $13.1{\pm}7.8$
w-r $6.4{\pm}0.1$ $4.2{\pm}1.4$ $1.5{\pm}0.7$ $3.4{\pm}2.0$ $0.0{\pm}0$ 0.0 $\textbf{20.2}{\pm}\textbf{2.9}$
half-m $47.7{\pm}0.2$ $54.9{\pm}0.2$ $48.1{\pm}0.1$ $50.0{\pm}0.2$ $\textbf{64.8}{\pm}\textbf{0.8}$ $54.0{\pm}0.0$
hp-m $55.4{\pm}2.9$ $42.7{\pm}37.6$ $57.4{\pm}2.3$ $74.3{\pm}6.0$ $\textbf{100.6}{\pm}\textbf{1.0}$ $97.9{\pm}4.0$
w-m $77.5\pm 2.3$ $73.6\pm 2.0$ $83.7\pm 1.2$ $86.3\pm 0.4$ $\textbf{93.8}\pm\textbf{1.0}$ $84.6\pm 0.3$
half-m-r $43.7\pm 0.3$ $51.8\pm 0.0$ $45.1\pm 9.8$ $46.9\pm 0.6$ $\textbf{53.6}\pm\textbf{1.0}$ $50.1\pm 0.4$
hp-m-r $55.4\pm 0.4$ $94.4\pm 4.0$ $62.7\pm 20.9$ $\textbf{100.7}\pm\textbf{0.5}$ $99.8\pm 0.8$ $98.8\pm 0.6$
w-m-r $69.3\pm 5.8$ $84.0\pm 3.6$ $76.3\pm 4.1$ $\textbf{93.8}\pm\textbf{0.5}$ $86.3\pm 1.6$ $\textbf{93.7}\pm\textbf{2.7}$
half-m-e $86.7\pm 4.4$ $80.4\pm 12.7$ $93.2\pm 1.9$ $\textbf{96.0}\pm\textbf{0.1}$ $92.5\pm 0.7$ $\textbf{96.1}\pm\textbf{0.6}$
hp-m-e $44.2\pm 4.3$ $99.2\pm 6.1$ $96.8\pm 3.8$ $107.6\pm 4.0$ $\textbf{109.3}\pm\textbf{5.2}$ $105.8\pm 0.9$
w-m-e $109.3\pm 1.0$ $101.2\pm 17.6$ $110.2\pm 0.0$ $109.9\pm 0.2$ $\textbf{115.2}\pm\textbf{1.7}$ $109.8\pm 0.0$
Average $51.2\pm 2.1$ $60.9\pm 8.4$ $58.0\pm 4.4$ $66.4\pm 1.2$ $\textbf{72.7}\pm\textbf{1.2}$ $70.3\pm 1.8$
@@ 5 Experiment
The objective of the experiments presented in this section is to validate the hypothesis that the utilization of the hypercube policy regularization framework can enhance the performance of policy regularisation algorithms in offline RL. To this end, our algorithm was validated using the D4RL dataset, a popular benchmark for this field, and the results were compared with those obtained by other well-known algorithms.
A variety of tasks were considered, drawn from a range of domains within the D4RL benchmark, including Gym, AntMaze and Adroit.
@@ 5.1.1 Experiment setting
Gym datasets are currently the most commonly used evaluation criterion tasks. Gym datasets typically have high-quality training curves, and the reward functions in Gym datasets are relatively smooth. Therefore, Gym datasets are relatively simple compared with other tasks.
To verify the efficacy of the hypercube policy regularization framework, it was integrated with TD3-BC and Diffusion-QL to create the TD3-BC-C and Diffusion-QL-C algorithms. The algorithms were then tested on the D4RL dataset, and the test results were compared with some classical and advanced algorithms. The compared algorithms include TD3-BC[7], which is one of the benchmark algorithms; the Diffusion-QL algorithm[28], which is the sota policy constraint algorithm and constitutes another benchmark algorithm; IQL[15], which is a policy constraint algorithm that uses a combination of V-value and Q-value functions; CQL[17], a classical Q-value constraint algorithm.
Training is performed for 2000 epochs in the medium-expert environment and 1000 epochs for other types of environments, 5 random seeds are used for each environment. The results of the training can be seen in Table LABEL:result, and the training curves can be seen in FigureLABEL:exp.
@@ 5.1.2 Experiment result
        maze-m-d         $3.2{\pm}2.5$         $0.0{\pm}0$ 0.0         $\textbf{62.1}{\pm}\textbf{5.1}$         $0.0{\pm}0$ 0.0
        maze-m-p         $0.4{\pm}0.6$         $0.0{\pm}0$ 0.0         $\textbf{59.9}{\pm}\textbf{8.7}$         $0.4{\pm}3.5$
        maze-l-d         $0.0{\pm}0$ 0.0         $14.2{\pm}29.2$         $25.2{\pm}6.7$         $\textbf{31.6}{\pm}\textbf{32.6}$
        maze-l-p         $0.0\pm 0$ 0.0         $2.2\pm 5.6$         $20.1\pm 6.7$         $\textbf{25.5}\pm\textbf{14.9}$
@@
Table LABEL:result demonstrates that TD3-BC-C exhibits optimal performance in $7$ out of $12$ environments. When compared with TD3-BC, it shows better performance in all environments except walked-random. Moreover, the performance of TD3-BC-C is superior to that of the current state-of-the-art algorithm, Diffusion-QL. Diffusion-QL-C displays a markedly enhanced performance in environments of inferior quality such as random environments. Furthermore, the overall performance of Diffusion-QL-C is also improved compared with Diffusion-QL.
It can be seen that although the performance of Diffusion-QL is much better than TD3-BC, the performance of TD3-BC-C is better than Diffusion-QL-C. It is hypothesized that this discrepancy can be attributed to the fact that Diffusion-QL has more hyperparameters compared with TD3-BC. The use of excessive hyperparameters may result in a reduction in the sensitivity of the hypercube policy regularization framework, thereby limiting the efficacy of the hypercube policy regularization framework in Diffusion-QL. Nevertheless, it is hypothesized that Diffusion-QL-C may demonstrate remarkable performance if the hyperparameters are subjected to further adjustment.
5.2 Experiment in other D4RL environment
To further validate the performance of the algorithms, additional experiments were conducted in Adroit and Maze environments. Compared with Gym environments, Adroit and Maze environments are more complex and challenging for agents to learn. Maze environments comprise a series of maze navigation tasks, wherein the agent is required to develop a policy to reach the end point from the path information provided in the static dataset. The agent is rewarded solely upon reaching the designated endpoint. It is therefore necessary for the algorithm to possess the ability to overcome sparse rewards. Adroit environments consist of some datasets that simulate a range of human activities. The complexity of human activities, the limited scope of the data and the relatively sparse rewards make it extremely challenging to learn Adroit environments.
The original reward was utilized for the experiments conducted on Gym and Adroit. For the experiments on Maze, since the majority of the algorithms were unable to achieve effective results with the original reward information. Consequently, a modification was made to this information following algorithms such as Diffusion-QL. In the Maze and Adroit environments, the algorithms employed for comparison are TD3-BC and Diffusion-QL, as the special parameters used by most algorithms in these environments have not been made available for public scrutiny.
@@ 5.2 Experiment in other D4RL environment
In order to ensure a fair comparison, the results presented in Table LABEL:result2 include the results obtained after training for 1000 epochs as well as the highest score achieved during the training process. This is due to the significant performance fluctuations of algorithms in maze environments. The training diagrams are illustrated in FigureLABEL:exp2.
The experimental results demonstrate that TD3-BC-C and Diffusion-C are also effective in complex environments, which further demonstrates the effectiveness of the hypercube policy regularization framework.
5.3 Hyperparameter setting
Table 3: The number of the state space and action space has been partitioned in different D4RL environments and the uncertainty constraint parameter $\delta$.
Environment state_dim $\delta$ epoch
hp-r 3 20 1000
hp-m-r 3 20 1000
hp-m 3 20 1000
hp-m-e 3 20 2000
half-r 6 5 1000
half-m-r 6 5 1000
half-m 6 5 1000
half-m-e 6 5 2000
w-r 6 5 1000
w-m-r 6 5 1000
w-m 6 5 1000
w-m-e 6 5 2000
maze-d 2 5 1000
maze-m 2 5 1000
maze-l 2 5 1000
pen-c 24 5 1000
pen-h 24 5 1000
The hypercube constraint framework introduces only one additional hyperparameter $\delta$, which represents the segmentation accuracy of the state space. Therefore, the value of $\delta$ can be determined based on its practical significance. In particular, the hyperparameter $\delta$ can be selected based on the dimensionality of the state space and the quality of the dataset.
For low-quality datasets, a lower $\delta$ is selected to permit the agent to explore more actions. In contrast, for higher-quality datasets, a higher $\delta$ is chosen, whereas for static datasets, policy cloning is deemed sufficient. As for the dimension of the space, the probability of two state actions occurring in the same hypercube is lower for higher dimensional spaces when $\delta$ remains constant. Consequently, a lower $\delta$ is selected for higher dimensional environments and a higher $\delta$ is selected for lower dimensional environments. In order to ensure a fair comparison, the majority of environments in this article have been configured with identical parameters. The selection of specific parameters in complex Maze and Adroit environments may result in better performance.
Table 3 illustrates the specific $\delta$ that was employed and lists some of the parameters that were deemed to be relevant.
5.4 Computational cost
Table 4: The computational cost of some algorithms in hopper medium. GPU Memory is the average of GPU Memory used in hopper-medium.
Algorithm Runtime(s) GPU Memory(GB)
IQL 14.6 2.0
CQL 62.2 1.4
TD3-BC 13.8 1.9
Diffusion-QL 68.9 2.0
TD3-BC-C 21.0 1.9
Diffusion-QL-C 77.3 2.0
To demonstrate the efficacy of the hypercube policy regularization framework from a different perspective, the runtime and space requirements of each algorithm are compared. The running time and space were obtained by executing hopper-medium for one epoch, the results are presented in Table 4. Table 4 reveals that TD3-BC-C and Diffusion-QL-C only increase a small amount of runtime and GPU Memory compared with TD3-BC and Diffusion-QL. It can therefore be concluded that the hypercube policy regularization framework has significant advantages in terms of performance and computational cost.
@@ 6 Conclusions
In this paper, the hypercube policy regularization framework is proposed, a novel framework for policy regularisation. This framework alleviates the constraints of policy regularization methods by enabling the agent to explore actions corresponding to similar states. It is indicated both theoretically and experimentally that the hypercube constraint framework can improve the performance of algorithms with policy regularization. This is achieved by alleviating constraints on OOD state actions and enabling the agent to explore some actions. In D4RL experiments, TD3-BC-C and Diffusion-QL-C obtained by the hypercube policy regularization framework demonstrated favorable results with the introduction of only one additional hyperparameter and a modest increase in computational cost. The performance of the hypercube policy regularization framework demonstrates that further research on the optimal utilization of static datasets has important implications for offline reinforcement learning. Therefore, we intend to pursue further research into the utilization of static datasets.
Appendix A Ablation experiment
As for the sensitivity of the parameter $\delta$, we illustrate that the hyperparameter $\delta$ has a negligible impact on the performance of the algorithm. Experiments are conducted with different $\delta$ in walker2d-medium and hopper-medium-replay environments for testing. The experimental results are shown in Figure LABEL:abl.
TD3-BC-C and Diffusion-QL-C are equivalent to TD3-BC and Diffusion-QL respectively When the value of a is very large. Therefore, TD3-BC and Diffusion-QL are employed to replace the situation when $\delta$ selects a larger value in the ablation experiment.
It can be seen from Figure LABEL:abl, in walker2d-medium-replay and hopper-medium, TD3-BC-C and Diffusion QL-C can both have results when the value of $\delta$ is not too extreme. However, when $\delta=1$, the exploration range of the agent is excessive, and the constraints on the out-of-distribution state actions are insufficient, rendering unsatisfactory results in some environments. When $\delta$ is excessively large, the hypercube policy regularization framework merely clones the behavior of the current state action, yielding only suboptimal outcomes.
Appendix B Parameters
Table 1: Partial hyperparameters in GPC-SAC
Hyperparameter value
discount 0.99
policy_rl 3e-4
qf_rl 3e-4
soft_target_tau 5e-3
train_pre_loop 1000
Q-network FC(256,256,256)
Optimizer Adam
activation_function relu
$\beta$ 1
$\beta_{next}$ 0.1
$\phi$ 2.5,40.0
$\eta$ 1.0
For the hyperparameters utilized in the Gym experiment, since our algorithm has been modified from TD3-BC and Diffusion-QL. Consequently, the majority of our hyperparameters are consistent with those of these two algorithms, as illustrated in Table 1.
In Adroit and Maze environments, refer to the parameter in  [28, 21]. The selection of parameter $\phi$ is $\{0.1,2.5,10,25\}$ for {pen,maze-u,maze-m,maze-l}, parameter $\eta$ as $\{0.1,0.15,0.5,2.0,2.0,3.0,4.5,3.5\}$ for {pen-c,pen-h,maze-u,maze-u-d,maze-m-p,maze-m-d,maze-l-p,maze-l-d}.
Appendix C Implement Detail
Our experimental results were obtained through the following method. The specific implementation of TD3-BC-C, Diffusion-B can be found at https://github.com/lastTarnished/GPC-SAC. Experiments are conducted on D4RL, which can be publicly used at https://github.com/Farama-Foundation/D4RL. The following are the implementations of the comparative algorithms in our experiment. (1)For CQL, https://github.com/lastTarnished/Hypyercud-Policy-Regularization/ has been referenced, which is the official implementation of CQL. (2)For the implementation of IQL, https://github.com/ikostrikov/implicit_q_learning has been referenced, this code has simplified the implementation of IQL compared with the official code, resulting in a slight decrease in performance in a few environments. (3)For the implementation of TD3-BC, https://github.com/sfujim/TD3_BC and no changes were made to the core code. (4)For the implementation of Diffusion-QL, https://github.com/Zhendong-Wang/Diffusion-Policies-for-Offline-RL has been referenced and no changes were made to the core code.

### v2

@@ 3.3 Algorithms
Input: Initial policy parameter $\it\theta$}; min-batch size $k$; Q-function parameters $\phi_{1}$$\phi_{2}$; Target Q-function parameters $\phi_{1}^{-}$, $\phi_{2}^{-}$; offline dataset $\it D$}; actions obtained by hypercube policy regularization $a_{\rm max}$
@@
1:for epoch=1 to MaxEpoch do
2:  for step=1 to MaxStep do
3:   Sample $k$ state-action pairs $\left\{(s_{t},a_{t}),t=1,..,k\right\}$ from D{D}
4:   Sample $a_{t+1}\sim\pi_{\theta}(a_{t+1}|s_{t+1})$ accroding to policy
5:   Update $Q_{{\phi_{1}}}$ and $Q_{{\phi_{2}}}$ by : $\begin{array}[]{l}E_{({s_{t}},a_{t},s_{t+1})\sim D,a_{t+1}\sim{\pi_{\theta}}}[(r({s_{t}},{a_{t}})+\gamma\min Q_{{\phi_{i}}}({s_{t+1}},{a_{t+1}})\\ -Q_{\phi_{i}}({s_{t}},{a_{t}}))^{2}]\end{array}$
6:   Compared $\mathop{\min}\limits_{i=1,2}{Q_{{\phi_{i}}}}(s,a)$ and $\mathop{\min}\limits_{i=1,2}{Q_{{\phi_{i}}}}(s,a_{\rm max})$ to Update $a^{\prime}$}
7:   Update policy by (14)
8:   Update Q target by: $\phi_{i}^{-}=\rho\phi_{i}^{-}+(1-\rho)\phi_{i}$
9:  end for
10:end for
This section outlines how the hypercube policy regularization framework can be integrated with policy regularization algorithms like TD3-BC and Diffusion-QL.
Taking TD3-BC as an example, our algorithm made only two modifications to TD3-BC. Firstly, a hypercube was constructed using Equation(2) and employed to store the optimal action $a_{\rm max}$within the hypercube. Each time the state $s$ corresponding to the hypercube is selected, a comparison is made between $Q(s,a)$ and $Q(s,a_{\rm max})$ to determine whether action $a$ should be replaced within the hypercube. Secondly, in policy updates, the action $a_{\rm max}$ from the hypercube is employed in place of $a$ from the static dataset for policy regularization. The policy update equation is given by:
$\arg\min L(\theta)={L_{c}}(\theta)-{L_{q}}(\theta)$ (14)
Where $\theta$ is the policy parameter, and $L_{c}$ is the policy regularization loss function.
@@
$L_{c}=E_{(s,a)\sim D}[{(a-\pi_{\theta}(s))^{2}}]$ (15)
Unlike the $L_{c}$ above, in our method, the $L_{c}$ is:
$L_{c}=E_{(s,a)\sim D}[{(a_{\rm max}-\pi_{\theta}(s))^{2}}]$ (16)
Given that $Q(s,a_{\rm max})\geq Q(s,a)$ and Theorem 3.1, it follows that the actions $a_{\rm max}$ within the hypercube will produce results that are either equal to or better than those produced by $a$ within the static dataset. Therefore, the utilization of (2) for behavior cloning can lead to a more effective policy.
@@
$L_{q}=\phi E_{s_{i}\sim D,a_{i}^{\prime}\sim\pi_{\theta}}[\sum_{i=0}^{T-1}Q(s_{i},a_{i}^{\prime})]/E_{(s_{i},a_{i})\sim D}[\sum_{i=0}^{T-1}\left|Q(s_{i},a_{i})\right|]$ (17)
@@
Here $\gamma=\eta/{E_{(s_{i},a_{i})\sim D}}[\sum_{i=0}^{T-1}|{Q_{\phi}}(s_{i},a_{i})|]$, $\eta$ is a hyperparameter.
@@ 4 Experiment
In this section, the hypercube policy regularization framework is validated through comprehensive evaluations of TD3-BC-C and Diffusion-QL-C across multiple D4RL environments. All of our code is open source [^1] https://github.com/lastTarnished/Hypercube-Policy-Regularization.
@@ 4.1.1 Experiment setting
Gym datasets are currently the most commonly used evaluation criterion tasks. Gym datasets typically have high-quality training curves, and the reward functions in Gym datasets are relatively smooth.
To verify the efficacy of the hypercube policy regularization framework, TD3-BC-C and Diffusion-QL-C were tested, and the test results were compared with some classical and advanced algorithms. The compared algorithms include TD3-BC[5], which is one of the benchmark algorithms; the Diffusion-QL algorithm[20], which is the sota policy regularization algorithm and constitutes another benchmark algorithm; IQL[10], which is a policy regularization algorithm that uses a combination of V-value and Q-value functions; CQL[12], a classical Q-value regularization algorithm.
Training is performed for 2000 epochs in the medium-expert environment and 1000 epochs for other types of environments, 5 random seeds are used for each environment. The results of the training are presented in Table LABEL:result.
@@ 4.1.2 Experiment result
Table LABEL:result demonstrates that TD3-BC-C exhibits optimal performance in $7$ of $12$ benchmark environments. When compared with TD3-BC, it shows better performance in all environments except walked-random. Moreover, in suboptimal static datasets (random/medium environments), the performance of TD3-BC-C and Diffusion-QL-C demonstrating 30% and 20% improvements over the SOTA algorithm Diffusion-QL, with overall performance enhancements of 9% and 5% across benchmark tests. These results demonstrate the effectiveness of the proposed method in enhancing the exploration capabilities of policy-constrained algorithms, particularly in suboptimal static datasets.
Notwithstanding these gains, in random environments, while TD3-BC-C and Diffusion-QL-C algorithms demonstrate superior performance, their performance in the lowest-quality datasets remains surpassed by SOTA algorithms[3]. Consequently, room for improvement persists in such algorithms for low-quality static datasets.
Env IQL CQL TD3-BC Diffusion TD3-BC-C Diffusion-C
half-r $11.0{\pm}3$ $\textbf{28.3}{\pm}\textbf{1}$ $11.7{\pm}1$ $21.0{\pm}1$ $26.5{\pm}2$ $23.8{\pm}1$
hp-r $7.8{\pm}0$ $16.4{\pm}15$ $8.5{\pm}0$ $8.1{\pm}0$ $\textbf{31.3}{\pm}\textbf{0.0}$ $13.1{\pm}8$
w-r $6.4{\pm}0$ $4.2{\pm}1$ $1.5{\pm}1$ $3.4{\pm}2$ $0.0{\pm}0$ $\textbf{20.2}{\pm}\textbf{3}$
half-m $47.7{\pm}0$ $54.9{\pm}0$ $48.1{\pm}0$ $50.0{\pm}0$ $\textbf{64.8}{\pm}\textbf{1}$ $54.0{\pm}0$
hp-m $66.2{\pm}3$ $42.7{\pm}38$ $57.4{\pm}2$ $74.3{\pm}6$ $\textbf{100.6}{\pm}\textbf{1}$ $97.9{\pm}4$
w-m $77.5\pm 2$ $73.6\pm 2$ $83.7\pm 1$ $86.3\pm 0$ $\textbf{93.8}\pm\textbf{1}$ $84.6\pm 0$
half-m-r $43.7\pm 0$ $51.8\pm 0$ $45.1\pm 10$ $46.9\pm 1$ $\textbf{53.6}\pm\textbf{1}$ $50.1\pm 0$
hp-m-r $55.4\pm 0$ $94.4\pm 4$ $62.7\pm 21$ $\textbf{100.7}\pm\textbf{1}$ $99.8\pm 1$ $98.8\pm 1$
w-m-r $69.3\pm 6$ $84.0\pm 4$ $76.3\pm 4$ $\textbf{93.8}\pm\textbf{1}$ $86.3\pm 2$ $\textbf{93.7}\pm\textbf{3}$
half-m-e $86.7\pm 4$ $80.4\pm 13$ $93.2\pm 2$ $\textbf{96.0}\pm\textbf{0}$ $92.5\pm 1$ $\textbf{96.1}\pm\textbf{1}$
hp-m-e $1.5\pm 4 99$$9.2\pm 6$ $96.8\pm 4$ $107.6\pm 4$ $\textbf{109.3}\pm\textbf{5}$ $105.8\pm 1$
[…]
Average $56.0\pm 2$ $60.9\pm 8$ $58.0\pm 4$ $66.4\pm 1$ $\textbf{72.7}\pm\textbf{1}$ $70.3\pm 2$
4.2 Experiment in other D4RL environment
To further validate the performance of the algorithms, additional experiments were conducted in Adroit and Maze environments. Maze environments comprise a series of maze navigation tasks. The agent is rewarded solely upon reaching the designated endpoint. It is therefore necessary for the algorithm to overcome sparse rewards. Adroit environments consist of some datasets that simulate a range of human activities. The complexity of human activities, the limited scope of the data and the relatively sparse rewards make it extremely challenging.
@@ 4.2 Experiment in other D4RL environment
maze-m-d $3.2{\pm}2.5$ $0.0{\pm}0.0$ $\textbf{62.1}{\pm}\textbf{5.1}$ $0.0{\pm}0.0$
maze-m-p $4{\pm}0.6 0.0$$0{\pm}0.0$ $\textbf{59.9}{\pm}\textbf{8.7}$ $0.4{\pm}3.5$
maze-l-d $0.0{\pm}0.0$ $14.2{\pm}29.2$ $25.2{\pm}6.7$ $\textbf{31.6}{\pm}\textbf{32.6}$
maze-l-p $0.0\pm 0.0$ $2.2\pm 5.6$ $20.1\pm 6.7$ $\textbf{25.5}\pm\textbf{14.9}$
@@
To ensure a fair comparison, Table LABEL:result2 includes the score after training for 1000 epochs and the highest score during the training process. This is due to the significant performance fluctuations in Maze environments. More experiment results are presented in the Appendices [^2] https://arxiv.org/abs/2411.04534.
In the Maze environment experiments, satisfactory performance was not achieved by either baseline algorithm. This outcome is attributed to the environment’s requirement for extensive exploration to identify optimal navigation paths, which is not met by the conservative nature of Q-constrained algorithms. On the other hand, experimental results in the Adroit environment, which necessitates meticulous imitation of complex human motions, demonstrate that the hypercube policy regularization framework does not compromise imitation capability. Rather, it enables refinement of imitated actions, leading to enhanced performance.
@@ 5 Conclusions
To resolve the excessive constraint issue in policy regularization methods that restricts further policy improvement, a hypercube policy constraint framework is proposed. The framework expands the policy optimization space while maintaining constraint strength through a hypercube-based partitioning mechanism. By integrating this framework with baseline algorithms TD3-BC and Diffusion-QL, the TD3-BC-C and Diffusion-QL-C algorithms are developed. Experimental results on the D4RL dataset demonstrate that TD3-BC-C and Diffusion-QL-C achieve SOTA performance among policy-constrained algorithms. Due to its ability to reduce the dependency of policy constraint algorithms on dataset quality, a promising research direction is established for future research.
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China [grant numbers: 12201656], Science and Technology Projects in Guangzhou [grant numbers: SL2024A04J01579] and Key Laboratory of Information Systems Engineering (CN).